NH2 Η trace H2SO4 -NH
Answer #1

ЛИ, NA NH - (Hz acho > и и, со it is one condensation kind af RAN- III TH₂O

Je + M N — 2 — .. NH-2 t-tho Cut) >C=NH-2 J-H* Y =N-2 also Nrz-C-NH - NH NH 2 - love pair ist and and are low neeleophile but